OpenStudy (anonymous):

A circle is shown on the coordinate plane below. What is the equation of this circle? (x – 1)2 + y2 = 2 (x + 1)2 + y2 = 4 (x + 1)2 + y2 = 2 (x – 1)2 + y2 = 4

OpenStudy (anonymous):

WHERE IS CIRCLE?

OpenStudy (anonymous):

here is the circle

OpenStudy (anonymous):

ITS A

OpenStudy (anonymous):

thank you (:

OpenStudy (anonymous):

SO you are my fan

OpenStudy (anonymous):

yes lol

OpenStudy (anonymous):

lol????????????????

OpenStudy (phi):

There are two ways to answer this question. If you don't know the equation of the circle, you could take a point on the circle (say x=-1, y=0), and try these in each of the possible answers. The correct equation will be true. or, you might remember that the equation of a circle centered at (0,0) is \[ x^2+y^2= r^2 \] where r is the radius. Find the diameter (side to side, or bottom to top) and divide by 2 Here the diameter is 4, so the radius is 4/2= 2. and r^2= 2*2=4 so look for an equation with 4 on the right side. Then use the first trick to pick the correct answer. or you might know the equation of the a circle centered at (x0,y0) is \[ (x-x_0)^2 + (y-y__0)^2 =r^2 \] to find the center, find the average of the x values on each side (-1+3)/2= 2/2= 1 for x and y= (-2+2)/2=0 so x0= 1 and y0=0 use these in the equation to get the answer.

OpenStudy (phi):

btw, the answer is not A